# Exploiting separation-dependent coherence to boost optical resolution

### v1

@@ I Introduction
In the second example, we resolve the separation between two interacting dipole emitters. We show that the dependence of the emission characteristics on the separation can drastically boost the sensitivity of the estimation. As a result, one can obtain useful information about the dipoles’ separation on a timescale significantly larger than the radiative lifetime of the excited state of the dipole. For the late stages of decay, even though the probability of detecting a photon is very small, the detection event contains a lot of information because of the strong interaction-induced correlations between the emitters.
@@ IV Measurement sensitivity
To describe the sensitivity of the SPADE measurement, we use the method of moments [14]. It sets the bound for estimators $\tilde{d}(N_{m})$ that are based on the mean values of the observables $N_{m}=\langle\hat{N}_{m}\rangle$:
@@
where $\mu$ is the number of measurement repetitions (assumed to be large) and the sensitivity $M$ is defined as
$M=\sum_{m,n=0}^{K}\frac{\partial N_{m}}{\partial d}\left(\mathbf{\Gamma}^{-1}\right)_{mn}\frac{\partial N_{n}}{\partial d},$ (14)
with $\mathbf{\Gamma}$ being the covariance matrix of the measured observables, with elements
$\Gamma_{mn}=\langle\hat{N}_{m}\hat{N}_{n}\rangle-\langle\hat{N}_{m}\rangle\langle\hat{N}_{n}\rangle.$ (15)
Moreover, the method of moments presents an optimal estimator [14]. This estimator relies on the linear combination of mean values $N_{m}$, and asymptotically saturates the bound given by (13) for a large enough number of measurement repetitions $\mu$.
@@ V Constant mutual coherence
V.1 Real-valued coherence ($\boldmath{\gamma=\gamma_{0}}$).
@@
In this scenario, the sensitivity expressed in (24) coincides with the quantum Fisher information [10, 11]. This implies the possibility of constructing an estimator $\tilde{d}(N_{m})$ based solely on a linear combination of the mean photon count values in any basis with the parity defined in Eq. (5) and achieving the ultimate resolution limit dictated by the quantum Cramér-Rao bound.
@@
V.2 Imaginary coherence ( $\boldmath{\gamma=i\gamma_{0}}$).
@@ VI.2 Interacting emitters
$\frac{d}{dt}\hat{\rho}=-if_{12}[\hat{\sigma}_{1}^{+}\hat{\sigma}_{2}^{-}+h.c.,\hat{\rho}]+\\ \frac{1}{2}\sum_{j,l=1}^{2}g_{jl}\left(2\hat{\sigma}_{j}^{-}\hat{\rho}\hat{\sigma}_{l}^{+}-\hat{\sigma}_{j}^{+}\hat{\sigma}_{l}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{j}^{+}\hat{\sigma}_{l}^{-}\right),$ (45)
where $\sigma^{\pm}_{j}$ are the dipoles’ transition operators. The first term in Eq. (45) is proportional to the collective frequency shift $f_{12}$ of the energy levels [30, 31]. The second, nonunitary term describes the spontaneous emission of the individual emitters ($j=l$) and their collective radiative decay ($j\neq l$). In the case of identical dipoles,
@@
where $\tau$ is the natural lifetime of the excited state of an individual dipole, and $z=2\pi d/\lambda$, with $\lambda$ the wavelength of the dipole transition.
@@
Its analytical form is presented in Appendix A. The coherency matrix of the excited modes remains unaffected by the collective shifts $f_{12}$. However, as discussed in Appendix A, the emission spectrum is influenced by $f_{12}$. Consequently, the measurement results and the sensitivity of the measurement are independent of $f_{12}$ exclusively in the considered case of frequency-blind measurements.
@@
Both the unitless emission rate $\dot{N}=N\tau/\Delta t=\Gamma^{(1)\hat{s}(t)}_{11}\tau/\Delta t$ and the degree of mutual coherence $\gamma=\Gamma^{(1)\hat{s}}_{12}/\Gamma^{(1)\hat{s}}_{11}$ depend on both time and separation, however, we will not explicitly indicate it in our notations. We display plots of these variables for various time moments $t$ as functions of the separation $d$ normalized with wavelength $\lambda$ in Fig. 7.
Figure 7: Properties of the dipoles radiation: the normalised emission rate $\dot{N}$ and the degree of mutual coherence $\gamma$ vs. separation $d$. The dashed line corresponds to the time $t=0.02\tau$, solid line $t=5\tau$.
One can see that at an early stage of the emission ($t=0.02\tau$) the dipoles fluoresce almost independently, with the emission rate close to that of an individual dipole, $\dot{N}\approx e^{-t/\tau}$, and with almost no coherence, $\gamma\approx 0$. However, after some time, the dipole-dipole interaction creates correlations and the difference with the individual dipole emission becomes evident in both the emission rate [34] and the mutual coherence. Crucially, this difference depends on the separation between the dipoles, which allows for a measurement of the separation with higher accuracy. The mutual coherence exhibits a particularly strong dependence on separation during the late emission stages ($t>\tau$) for separations $d\approx n\frac{\lambda}{2},n\in\mathbb{N}$. Therefore, we can expect heightened sensitivity in estimating separation around these points. Note that due to the geometrical symmetry of the problem the mutual coherence $\gamma$ stays real at any moment, i.e. $\gamma(t)\in\mathbb{R}~{}\forall t$. However, all the conclusions below can be generalized to the case of the complex degree of mutual coherence by replacing $\gamma_{0}$ with $\operatorname{Re}~{}\gamma$.
The mean number of photons emitted by each dipole during a short time interval $\Delta t\ll\tau$ is low ($N\ll 1$). Furthermore, the imaging system is characterized by high losses, $\kappa\ll 1$. Therefore, irrespectively of the statistics of the sources, the detection statistics can be described by a Poisson distribution which also approximates that of faint thermal sources discussed earlier. Besides, the radiation emitted by the two dipoles in the far field represents a superposition of two spherical waves and therefore is locally equivalent to the radiation emitted by two-point sources (see Appendix A). Thus, in the paraxial approximation, the model developed in the previous sections can be applied to the problem of resolving dipoles.
@@
Figure 8: Normalized sensitivity rate $\dot{\mathcal{M}}$ (51) of two dipoles’ separation estimation vs. time $t$. The red line corresponds to $\sigma_{PSF}=\lambda$, the blue line to $\sigma_{PSF}=5\lambda$, and the dashed line to non-interacting dipoles.
Figure 9: Normalized sensitivity rate $\dot{\mathcal{M}}$ (51) of two dipoles’ separation estimation vs. separating $d$. The red line corresponds to $\sigma_{PSF}=\lambda$, the blue line to $\sigma_{PSF}=5\lambda$, and the dashed line to non-interacting dipoles.
@@
In Fig. 8 and 9, one can observe a pronounced enhancement in sensitivity within regions where the emission characteristics (the mutual coherence and the emission rate) exhibit rapid changes with a change in separation. Meanwhile, the sensitivity remains low around the extrema of the emission characteristics. In typical cases, the sensitivity for interacting dipoles is significantly larger and decaying significantly slower compared to non-interacting dipoles. Even after more than 15 lifetimes of the excited state, when the probability of photon detection is very low, the average sensitivity rate remains notably high due to the substantial amount of information in “late” detection events.
@@
The result of this calculation is shown in Fig. 10. One can see that the interaction effects can increase the total sensitivity of the separation estimation by several orders of magnitude around the points $d\approx n\frac{\lambda}{2},n\in\mathbb{N}$, where the mutual coherence of the emission strongly depends on the separation. In between these points, for the separations around $d\approx\left(n-\frac{1}{2}\right)\frac{\lambda}{2},n\in\mathbb{N}$, both the emission rate and mutual coherence exhibit extrema. Consequently, the sensitivity is not heightened by the separation-dependent emission characteristics around these specific points.
@@
Figure 10: Normalized total sensitivity $\mathcal{M}_{Tot}$  (52) of two dipoles’ separation estimation vs. separation $d$. The red line corresponds to $\sigma_{PSF}=\lambda$, the blue line to $\sigma_{PSF}=5\lambda$, and the dashed line to non-interacting dipoles.
@@
Curiously, the sensitivity boost is more pronounced for smaller separations, where the interaction between the emitters is stronger. As a result, in certain instances, one can achieve more precise estimations for smaller separations compared to larger ones.
@@ VII Conclusion
We have analyzed the problem of resolving partially coherent thermal sources with SPADE measurement. To estimate the sensitivity of a separation estimation, we utilized the method of moments, which allowed us to make no assumptions about the brightness of the sources. We have found analytical expressions for the sensitivity including the cases of separation-dependent mutual coherence and emission rate. We studied two specific examples of separation-dependent coherence: a reflection of light coming from a finite-coherence-width illumination source and creating mutual coherence due to the interaction of the emitters. In both cases, we demonstrate the possibility of a significant boost in the separation estimation sensitivity due to the additional mechanism of the parameter encoding to the problem. Our analysis shows that for efficient resolving of the reflecting objects one needs to use an illumination source with a narrow coherence width (on the order of the reflectors’ separation) and a short coherence time. This ensures the full advantage of the separation-dependent coherence of the reflected light. Examining the interacting emitters, we demonstrate that the sensitivity of separation estimation can be increased by several orders of magnitude compared to independent emitters. This enhancement arises from the separation-dependent mutual coherence and emission rates of the interacting dipoles.

### v3

@@ I Introduction
In the second example, we resolve the separation between two interacting dipole emitters. We show that the dependence of emission characteristics on the separation can drastically boost the estimation sensitivity. As a result, one can obtain useful information about the dipoles’ separation on a timescale significantly larger than the radiative lifetime of the excited state of the dipole. For the late stages of decay, even though the probability of detecting a photon is very small, the detection event contains a lot of information because of the strong interaction-induced correlations between the emitters. We show that this effect is resistant to the dephasing of the emitters and to the detection noise.
@@ IV Measurement sensitivity
To describe the sensitivity of the SPADE measurement, we use the method of moments [27, 14]. This method studies the estimators $\tilde{d}$ that are based only on the measured sample means $\overline{x}^{(\mu)}_{m}$ of the observables $\hat{X}_{m}$. The variance $\Delta^{2}\tilde{d}$ of such estimators is bounded as
@@
where $\mu$ is the number of measurement repetitions (assumed to be large, so that the central limit theorem can be applied and the statistics of the sample means $\overline{x}^{(\mu)}_{m}$ can be considered Gaussian). The sensitivity $M$ is defined as
$M=\sum_{m,n=0}^{K}\frac{\partial X_{m}}{\partial d}\left(\mathbf{\Gamma}^{-1}\right)_{mn}\frac{\partial X_{n}}{\partial d},$ (14)
with $X_{m}=\langle\hat{X}_{m}\rangle$ being the mean values of observables, and $\mathbf{\Gamma}$ the measurement covariance matrix, with elements
$\Gamma_{mn}=\langle\hat{X}_{m}\hat{X}_{n}\rangle-\langle\hat{X}_{m}\rangle\langle\hat{X}_{n}\rangle.$ (15)
The derivation of the bound (13) and a simple estimator that saturates this bound are presented in Appendix A.
In the considered scheme, the observables $\hat{X}_{m}$ are the photon number operators of the modes $f_{m}(x)$:
$\hat{X}_{m}=\hat{N}_{m}=\hat{a}^{\dagger}_{m}\hat{a}_{m}.$ (16)
@@ V Constant mutual coherence
V.1 Real-valued coherence ($\boldmath{\gamma=\gamma_{0}}$)
@@
In this scenario, the sensitivity expressed in (25) coincides with the quantum Fisher information [10, 11]. This implies the possibility of constructing an estimator $\tilde{d}(\{\overline{N}^{(\mu)}_{m}\})$ based solely on the observed sample mean numbers of counts $\overline{N}^{(\mu)}_{m}$ in any basis with the parity defined in Eq. (5) and achieving the ultimate resolution limit dictated by the quantum Cramér-Rao bound.
@@
V.2 Imaginary coherence ( $\boldmath{\gamma=i\gamma_{0}}$)
@@ VI.2 Interacting emitters
$\frac{d}{dt}\hat{\rho}=-if_{12}[\hat{\sigma}_{1}^{+}\hat{\sigma}_{2}^{-}+h.c.,\hat{\rho}]~{}+\\ \frac{1}{2}\sum_{j,l=1}^{2}g_{jl}\left(2\hat{\sigma}_{j}^{-}\hat{\rho}\hat{\sigma}_{l}^{+}-\hat{\sigma}_{j}^{+}\hat{\sigma}_{l}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{j}^{+}\hat{\sigma}_{l}^{-}\right)+\\ {\sum_{j=1}^{2}\eta_{j}\left(2\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}-\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}\hat{\rho}-\hat{\rho}\hat{\sigma}_{j}^{+}\hat{\sigma}_{j}^{-}\right)},$ (46)
where $\hat{\sigma}^{\pm}_{j}$ are the dipoles’ transition operators. The first term in Eq. (46) is proportional to the collective frequency shift $f_{12}$ of the energy levels [31, 32]. The second term describes the spontaneous emission of the individual emitters ($j=l$) and their collective radiative decay ($j\neq l$). In the case of identical dipoles,
@@
where $\tau$ is the natural lifetime of the excited state of an individual dipole, and $z=2\pi d/\lambda$, with $\lambda$ the wavelength of the dipole transition. The final term in the master equation accounts for the dephasing of the dipoles with the rates $\eta_{i}$ ($i=1,2$), due to their different local environments.
@@
The coherency matrix of the excited modes remains unaffected by the collective shifts $f_{12}$. However, as discussed in Appendix B, the emission spectrum is influenced by $f_{12}$. Consequently, the measurement results and the sensitivity of the measurement are independent of $f_{12}$ exclusively in the considered case of frequency-blind measurements.
@@
Both the unitless emission rate $\dot{N}=N\tau/\Delta t=\Gamma^{(1)\hat{s}(t)}_{11}\tau/\Delta t$ and the degree of mutual coherence $\gamma=\Gamma^{(1)\hat{s}}_{12}/\Gamma^{(1)\hat{s}}_{11}$ depend on both time and separation, however, we will not explicitly indicate it in our notations.
VI.2.1 Model without dephasing ($\eta_{1}=\eta_{2}=0$)
First, we examine the evolution of the dipoles’ state without taking into account the dephasing process. The analytical solution for the dipole correlators in this scenario is presented in Appendix B. The associated emission rate $\dot{N}$ and the degree of mutual coherence $\gamma$ for different time instances $t$ are depicted in Fig. 7 as functions of the separation $d$ normalized by the wavelength $\lambda$.
Figure 7: Properties of the dipoles radiation: the normalised emission rate $\dot{N}$ and the degree of mutual coherence $\gamma$ vs. separation $d$. The dashed line corresponds to the time $t=0.02\tau$, solid line $t=5\tau$. No dephasing ($\eta=0$).
One can see that at an early stage of the emission ($t=0.02\tau$) the dipoles fluoresce almost independently, with the emission rate close to that of an individual dipole, $\dot{N}\approx e^{-t/\tau}$, and with almost no coherence, $\gamma\approx 0$. However, after some time, the dipole-dipole interaction creates correlations and the difference with the individual dipole emission becomes evident in both the emission rate [35] and the mutual coherence. Crucially, this difference depends on the separation between the dipoles, which allows for a measurement of the separation with higher accuracy. Without accounting for the decoherence process, the mutual coherence exhibits a particularly strong dependence on separation during the late emission stages ($t>\tau$) for separations $d\approx n\frac{\lambda}{2},n\in\mathbb{N}$. Therefore, we can expect heightened sensitivity in estimating separation around these points. Note that due to the geometrical symmetry of the problem the mutual coherence $\gamma$ stays real at any moment, i.e. $\gamma(t)\in\mathbb{R}~{}\forall t$. However, all the conclusions below can be generalized to the case of the complex degree of mutual coherence by replacing $\gamma_{0}$ with $\operatorname{Re}~{}\gamma$.
The mean number of photons emitted by each dipole during a short time interval $\Delta t\ll\tau$ is low ($N\ll 1$). Furthermore, the imaging system is characterized by high losses, $\kappa\ll 1$. Therefore, irrespectively of the statistics of the sources, the detection statistics can be described by a Poisson distribution which also approximates that of faint thermal sources discussed earlier. Besides, the radiation emitted by the two dipoles in the far field represents a superposition of two spherical waves and therefore is locally equivalent to the radiation emitted by two-point sources (see Appendix B). Thus, in the paraxial approximation, the model developed in the previous sections can be applied to the problem of resolving dipoles.
@@ VI.2.1 Model without dephasing ($\eta_{1}=\eta_{2}=0$)
Figure 8: Normalized sensitivity rate $\dot{\mathcal{M}}$ (52) of two dipoles’ separation estimation vs. time $t$. The light blue line corresponds to $\sigma_{PSF}=5\lambda$, the red line to $\sigma_{PSF}=\lambda$, and the dashed line to non-interacting dipoles.
Figure 9: Normalized sensitivity rate $\dot{\mathcal{M}}$ (52) of two dipoles’ separation estimation vs. separation $d$. The light blue line corresponds to $\sigma_{PSF}=5\lambda$, the red line to $\sigma_{PSF}=\lambda$, and the dashed line to non-interacting dipoles. No dephasing ($\eta=0$).
@@
In Fig. 9, one can observe a pronounced enhancement in sensitivity within regions where the emission characteristics (the mutual coherence and the emission rate) exhibit rapid changes with a change in separation. Meanwhile, the sensitivity remains low around the extrema of the emission characteristics. In typical cases, the sensitivity for interacting dipoles is significantly larger and decaying significantly slower compared to non-interacting dipoles. Even after more than 15 lifetimes of the excited state, when the probability of photon detection is very low, the average sensitivity rate remains notably high due to the substantial amount of information in “late” detection events.
@@
where $\dot{\mathcal{M}}$ is defined in (52). The result of this calculation is shown in Fig. 10. One can see that the interaction effects can increase the total sensitivity of the separation estimation by several orders of magnitude around the points $d\approx n\frac{\lambda}{2},n\in\mathbb{N}$, where the mutual coherence of the emission strongly depends on the separation, if the dephasing effect is not present. In between these points, for the separations around $d\approx\left(n-\frac{1}{2}\right)\frac{\lambda}{2},n\in\mathbb{N}$, both the emission rate and mutual coherence exhibit extrema. Consequently, the sensitivity is not heightened by the separation-dependent emission characteristics around these specific points.
@@
Figure 10: Normalized total sensitivity $\mathcal{M}_{Tot}$  (53) of two dipoles’ separation estimation vs. separation $d$. The light blue line — $\sigma_{PSF}=5\lambda$, the red line — $\sigma_{PSF}=\lambda$, the dashed line — non-interacting dipoles. No dephasing ($\eta=0$).
VI.2.2 Model with dephasing ($\eta_{1},\,~{}\eta_{2}\neq 0$)
In realistic scenarios, the process of building up coherence through the dipole-dipole interactions competes with the coherence loss caused by dipoles’ individual dephasing. The master equation (46) with dephasing can be solved analytically, however the solution is too long and cumbersome to be presented within this paper. Nevertheless, it demonstrates that the coherency matrix for our system only depends on the average dephasing rate $\eta=(\eta_{1}+\eta_{2})/2$ of the dipoles. In the following, we discuss the impact of dephasing for different values of $\eta$. Figure 11 illustrates the emission characteristics (the emission rate and the mutual coherence) at the time instance $t=5\tau$. In the weak dephasing regime ($\eta=0.2/\tau$), where the dephasing time is substantially longer than the natural lifetime of the excited state $\tau$, the effect of dephasing has a limited influence on the emission process. On the other hand, when the dephasing process is faster than the emission ($\eta=2.0/\tau$), one can observe a significant decrease in the mutual coherence $\gamma$.
Figure 11: Emission characteristics: normalized emission rate $\dot{N}$ and degree of mutual coherence $\gamma$ at the time moment $t=5\tau$. The red line corresponds to the dephasing-free model $\eta=0$, the light blue line – weak dephasing $\eta=0.2/\tau$, the grey line – strong dephasing $\eta=2.0/\tau$.
As expected, the reduction in emission coherence due to the dephasing effect leads to a weaker sensitivity boost. Fig. 12 illustrates the total sensitivity of the separation estimation defined in (53). The top plot highlights the robustness of the considered scheme to weak dephasing, where the sensitivity is only minimally affected, compared to the dephasing-free case, shown in Fig. 10. Conversely, the bottom plot in Fig. 12 reveals that strong dephasing significantly diminishes the sensitivity boost arising from dipole interaction. However, even in this scenario, the sensitivity can be several times higher compared to non-interacting dipole emission.
Figure 12: Normalized total sensitivity $\mathcal{M}_{Tot}$  (53) for the model with dephasing. Top inset – weak dephasing $\eta=0.2/\tau$, bottom inset – strong dephasing $\eta=2.0/\tau$. The light blue line — $\sigma_{PSF}=5\lambda$, the red line — $\sigma_{PSF}=\lambda$, the dashed line — non-interacting dipoles.
VI.2.3 Model with detection noise
Another factor that can potentially significantly reduce sensitivity is the presence of detection noise. If the detection system produces dark counts or detects background light, then late rare detection events may be dominated by these false counts. To address this effect, we modify the detection covariance matrix (18) as
$\mathbf{\Gamma}_{\text{DC}}=\mathbf{\Gamma}+N_{\text{DC}}\mathbf{I},$ (54)
where $\mathbf{I}$ stands for the identity matrix, and $N_{\text{DC}}$ is the average number of dark counts per detection mode. Here, the dark counts are assumed to have Poisson statistics, being uncorrelated from each other and the measured signal.
Using the covariance matrix of noisy detection (54), we numerically calculate the sensitivity (14), taking into account the separation-dependent emission rate (65) and mutual coherence (66) of dipoles. We consider detection in the first four Hermite-Gauss modes and express the rate of dark counts relative to the rate of real counts at the early stages $\dot{N}_{0}=2\kappa$. Note, that the rate of real detection events exponentially decreases with time, while the rate of dark counts remains constant.
The sensitivity rate for noisy detection is presented in Fig. 13. Here we consider strong detection noise, with the dark count rate being 50% of the initial rate of real counts $\dot{N}_{\text{DC}}=0.5\dot{N}_{0}$. Comparing this plot to the ideal case (Fig. 8), one can observe a significant decrease in the sensitivity rate for late detection events. After $t=5\tau$ one obtains almost no information about the separation of dipoles. This decrease occurs because the dark count rate is much higher than the rate of real counts for these late stages. As a result, in this case, most of the useful information comes from the early and intermediate radiation stages ($t<5\tau$).
Figure 13: Normalized sensitivity rate $\dot{\mathcal{M}}$ of two dipoles’ separation estimation vs. time $t$ in case of noisy detection $\dot{N}_{\text{DC}}=0.5\dot{N}_{0}$.
Figure 14: Normalized total sensitivity $\mathcal{M}_{Tot}$ vs. separation $d$ for the model with detection noise. Top inset — low rate of the dark counts, bottom inset — intense noise.
Integrating the sensitivity rate over time we obtain the total sensitivity (53) for noisy detection and present it in Fig. 14. As expected, compared to the ideal case (Fig. 10), the sensitivity is suppressed in the presence of the detection noise. However, this is also the case for non-interacting dipoles. Thus, one can still see a significant increase in the sensitivity (two to three orders of magnitude for small separations) coming from the dipole interaction, even for relatively intense detection noise $\dot{N}_{\text{DC}}=0.5\dot{N}_{0}$.
@@ VI.2.3 Model with detection noise
Curiously, in all the considered cases the sensitivity boost is more pronounced for smaller separations, where the interaction between the emitters is stronger. As a result, in the context of single-parameter estimation, one can achieve higher precision in estimating smaller separations compared to larger ones.
@@ VII Conclusion
We have analyzed the problem of resolving partially coherent thermal sources with SPADE measurement. To estimate the sensitivity of a separation estimation, we utilized the method of moments, which allowed us to make no assumptions about the brightness of the sources. We have found analytical expressions for the sensitivity including the cases of separation-dependent mutual coherence and emission rate. We studied two specific examples of separation-dependent coherence: a reflection of light coming from a finite-coherence-width illumination source and creating mutual coherence due to the interaction of the emitters. In both cases, we demonstrate the possibility of a significant boost in the separation estimation sensitivity due to the additional mechanism of the parameter encoding to the problem. Our analysis shows that for efficient resolving of the reflecting objects one needs to use an illumination source with a narrow coherence width (on the order of the reflectors’ separation) and a short coherence time. This ensures the full advantage of the separation-dependent coherence of the reflected light. Examining the interacting emitters, we demonstrate that the sensitivity of separation estimation can be increased by several orders of magnitude compared to independent emitters. This enhancement arises from the separation-dependent mutual coherence and emission rates of the interacting dipoles. The effect remains robust in the presence of weak dephasing of dipoles. Although strong dephasing decreases the observed resolution boost, it does not eliminate it entirely. We have also demonstrated that the presence of detection noise reduces sensitivity but preserves the relative boost caused by the interaction of dipoles.